( ax/a+b ) = ( by/y+b ) = (cz/z+b ) & x, y, z, are in H.P. then a, b, c, are in


1) A.P.


2) G.P.


3) H.P.


4) None

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Mayank

( ax/x+b ) = ( by/y+b ) = (cz/z+b )

let

 ( ax/x+b ) = ( by/y+b ) = (cz/z+b ) =1/k

so x +b =axk   or    x =b/(ak-1)

 

   y+ b  =byk   or  y = b/(bk-1)

 

  z+b= czk   or   z =b/(ck-1)

 

 x ,y,z  are in H.P

or   1/x  ,1/y ,1/z   are in AP

or (ak-1)/b ,(bk-1)/b ,(ck-1)/b  are in AP

or (ak-1) ,(bk-1) ,(ck-1)  are in AP

or (ak) ,(bk1) ,(ck)  are in AP

or  a,b,c   are in AP

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin